# State-space kinetic Ising model reveals task-dependent entropy flow in sparsely active nonequilibrium neuronal dynamics

### v2

@@ Acknowledgments
This work was supported by JSPS KAKENHI Grant Number JP 20K11709, 21H05246, 24K21518, 25K03085.
Supporting Information
@@
Graduate School of Life Sciences, Hokkaido University, Sapporo, Japan
@@
1 S1 text. State-space kinetic Ising model
In this Supplementary Note, we provide the filtering and smoothing algorithms for the time-varying kinetic Ising models, and optimization method of its hyperparameters via Expectation Maximization algorithm.
@@ 1.1 Model
Let $x_{i,t}=\{0,1\}$ be an outcome of a binary random variable of neuron $i$ at time $t$ ($i=1,\ldots,N$, $t=0,\ldots,T$). In the kinetic Ising model, the activation of neuron $i$ at time $t$ independently depends on the activities of the neurons in the previous time step t1t-1. The probability mass function of $x_{i,t}$ is given as
$p({x}_{i,t}|{x}_{1,t-1},\ldots,{x}_{N,t-1},\boldsymbol{\theta}^{i}_{t})=\frac{\exp\left[\theta_{i,t}x_{i,t}+\sum_{j=1}^{N}\theta_{ij,t}x_{i,t}x_{j,t-1}\right]}{1+\exp\left[\theta_{i,t}+\sum_{j=1}^{N}\theta_{ij,t}x_{j,t-1}\right]},$ (76)
where $\theta_{i,t}$ is a time-dependent field parameter that determines the bias for inputs to the $i$-th neuron at time $t$, and $\theta_{ij,t}$ is a time-dependent coupling parameter from the $j$-th neuron to the $i$-th neuron. These parameters are collectively denoted as $\boldsymbol{\theta}^{i}_{t}=(\theta_{i,t},\theta_{i1,t},\ldots\theta_{ij,t},\ldots{\theta}_{iN,t})$. Using the log normalization function,
@@
$p({x}_{i,t}|{x}_{1,t-1},\ldots,{x}_{N,t-1},\boldsymbol{\theta}^{i}_{t})$ $=\exp\left[\theta_{i,t}x_{i,t}+\sum_{j=1}^{N}\theta_{ij,t}x_{i,t}x_{j,t-1}-\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}_{t-1})\right],$ (78)
The joint probability mass function is given as
$\prod_{i=1}^{N}p({x}_{i,t}|{x}_{1,t-1},\ldots,{x}_{N,t-1},\boldsymbol{\theta}^{i}_{t}).$ (79)
Typical neurophysiological experiments repeat multiple trials of measurement under the same experimental conditions. We let $x^{l}_{i,t}=\{0,1\}$ be a binary variable at time $t$ in the $l$-th trial ($i=1,\ldots,N$, $t=0,\ldots,T$, $l=1,\ldots,L$). We collectively denote the binary patterns of simultaneously recorded neurons at time $t$ in the $l$-th trial using a vector, $\mathbf{x}^{l}_{t}=(x^{l}_{1,t},\ldots,x^{l}_{N,t})$. Further, we denote the patterns at time step $t$ from all trials by $\mathbf{x}_{t}=(\mathbf{x}^{1}_{t},\ldots,\mathbf{x}^{l}_{t},\ldots,\mathbf{x}^{L}_{t})$ and denote all the patterns up to time $t$ by $\mathbf{x}_{0:t}$. Similarly, we use $\boldsymbol{\theta}_{t}=(\boldsymbol{\theta}^{1}_{t},\ldots,\boldsymbol{\theta}^{i}_{t},\ldots,\boldsymbol{\theta}^{N}_{t})$ and $\boldsymbol{\theta}_{1:t}$ for the time-varying parameters.
@@
$p(\mathbf{x}_{0:T}|\boldsymbol{\theta}_{1:T})$ $=\prod_{l=1}^{L}\prod_{i=1}^{N}\left[p(x^{l}_{i,0})\prod_{t=1}^{T}p({x}^{l}_{i,t}|\mathbf{x}^{l}_{t-1},\boldsymbol{\theta}^{i}_{t})\right],$ (80)
where we use $p(x^{l}_{i,0})=0.5$.
In the state-space model, the state model defines the discrete-time stochastic processes of the latent variables, which are the time-varying parameters $\boldsymbol{\theta}_{0:T}$ in our model. We use the following Gaussian model:
$p(\boldsymbol{\theta}_{0:T})$ $=\prod_{i=1}^{N}\left[p(\boldsymbol{\theta}^{i}_{0}|\boldsymbol{\mu}^{i},\mathbf{\Sigma}^{i})\prod_{t=1}^{T}p(\boldsymbol{\theta}^{i}_{t}|\boldsymbol{\theta}^{i}_{t-1},{Q}^{i})\right],$ (81)
@@
$p(\boldsymbol{\theta}^{i}_{t}|\boldsymbol{\theta}^{i}_{t-1},\mathbf{Q}^{i})$ $=\frac{1}{\sqrt{|2\pi\mathbf{Q}^{i}|}}\exp\left[\frac{1}{2}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})^{\top}(\mathbf{Q}^{i})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})\right]$ (82)
with $\mathbf{Q}^{\,i}$ being the noise covariance for the transition of the $i$-th neuron. The initial density of the $i$-th neuron $p(\boldsymbol{\theta}_{0}|\boldsymbol{\mu}^{i},\mathbf{\Sigma}^{i})$ is given as a Gaussian distribution with mean $\boldsymbol{\mu}^{i}$ and covariance $\mathbf{\Sigma}^{i}$. In practice, we used a zero vector and unit matrix, respectively.
The filter mean is given as
$\boldsymbol{\theta}^{i}_{t-1|t-1}=\int p(\boldsymbol{\theta}^{i}_{t-1}|\mathbf{x}_{0:t-1})\boldsymbol{\theta}^{i}_{t-1}d\boldsymbol{\theta}^{i}_{t-1}=E_{\boldsymbol{\theta}^{i}_{t-1}|\mathbf{x}_{0:t-1}}\boldsymbol{\theta}^{i}_{t-1}$ (83)
It represents the expected value of the parameter at time t1t-1 using data up to t1t-1. This is called the filter mean. Next, $\mathbf{W}^{i}_{t-1|t-1}$ is explained.
$\mathbf{W}^{i}_{t-1|t-1}=E_{\boldsymbol{\theta}^{i}_{t-1}|\mathbf{x}_{0:t-1}}[\boldsymbol{\theta}^{i}_{t-1}-E_{\boldsymbol{\theta}^{i}_{t-1}|\mathbf{x}_{0:t-1}}\boldsymbol{\theta}^{i}_{t-1}][\boldsymbol{\theta}^{i}_{t-1}-E_{\boldsymbol{\theta}^{i}_{t-1}|\mathbf{x}_{0:t-1}}\boldsymbol{\theta}^{i}_{t-1}]^{\top}$ (84)
That is, the covariance matrix of $\boldsymbol{\theta}_{t-1}$, which represents the deviation from the filter means above.
@@ 1.2 One-step prediction density
In this section, we derive the one-step prediction density $p(\boldsymbol{\theta}_{t}\mid\mathbf{x}_{0:t-1},\mathbf{w})$, using Chapman–Kolmogorov’s equation while utilizing completing the square.
We first specify the prior distribution at $t=1$ using a Gaussian distribution as
$p(\boldsymbol{\theta}_{1}\mid\mathbf{x}_{0},\mathbf{w})$ $=\mathcal{N}\bigl(\boldsymbol{\mu},\boldsymbol{\Sigma}\bigr),$ (85)
where $\boldsymbol{\mu}$ and $\boldsymbol{\Sigma}$ are the mean and covariance of this initial distribution.
For $t=2,\ldots,T$, the one-step prediction density $p(\boldsymbol{\theta}_{t}\mid\mathbf{x}_{0:t-1},\mathbf{w})$ is computed via the Chapman–Kolmogorov equation:
$p(\boldsymbol{\theta}_{t}\mid\mathbf{x}_{0:t-1},\mathbf{w})$ $=\int p(\boldsymbol{\theta}_{t},\boldsymbol{\theta}_{t-1}\mid\mathbf{x}_{0:t-1},\mathbf{w})\,d\boldsymbol{\theta}_{t-1}$
$=\int p(\boldsymbol{\theta}_{t}\mid\boldsymbol{\theta}_{t-1},\mathbf{x}_{0:t-1},\mathbf{w})\;p(\boldsymbol{\theta}_{t-1}\mid\mathbf{x}_{0:t-1},\mathbf{w})\,d\boldsymbol{\theta}_{t-1}.$ (86)
We assume that the filter density factors into a product of individual neurons. Coupled with the factorized assumption of the state model, this leads to the factorization of the one-step prediction density:
$p(\boldsymbol{\theta}_{t}\mid\mathbf{x}_{0:t-1},\mathbf{w})$ $=\prod_{i=1}^{N}\int p(\boldsymbol{\theta}^{\,i}_{t}\mid\boldsymbol{\theta}^{\,i}_{t-1},\mathbf{w})\;p(\boldsymbol{\theta}^{\,i}_{t-1}\mid\mathbf{x}_{0:t-1},\mathbf{w})\,d\boldsymbol{\theta}^{\,i}_{t-1}.$ (87)
We further assume that the filter distribution at time t1t-1, $p(\boldsymbol{\theta}^{\,i}_{t-1}\mid\mathbf{x}_{0:t-1},\mathbf{w})$, is approximated by a Gaussian distribution with mean $\boldsymbol{\theta}^{\,i}_{t-1\mid t-1}$ and covariance $\mathbf{W}^{\,i}_{t-1\mid t-1}$ (to be justified at the next filtering step):
$p(\boldsymbol{\theta}^{\,i}_{t-1}\mid\mathbf{x}_{0:t-1},\mathbf{w})=\mathcal{N}\bigl(\boldsymbol{\theta}^{\,i}_{t-1\mid t-1},\mathbf{W}^{\,i}_{t-1\mid t-1}\bigr).$ (88)
Given the Gaussian transition model $p(\boldsymbol{\theta}^{\,i}_{t}\mid\boldsymbol{\theta}^{\,i}_{t-1},\mathbf{w})=\mathcal{N}(\boldsymbol{\theta}^{\,i}_{t-1},\mathbf{Q}^{\,i})$, the one-step prediction density $p(\boldsymbol{\theta}_{t}\mid\mathbf{x}_{0:t-1},\mathbf{w})$ becomes a Gaussian distribution. Namely, by completing the square with resect to $\boldsymbol{\theta}^{\,i}_{t}$ and calculating the integral, we obtain
$p(\boldsymbol{\theta}_{t}\mid\mathbf{x}_{0:t-1},\mathbf{w})=\prod_{i=1}^{N}\mathcal{N}\bigl(\boldsymbol{\theta}^{\,i}_{t\mid t-1},\,\mathbf{W}^{\,i}_{t\mid t-1}\bigr),$ (89)
@@
$\boldsymbol{\theta}^{\,i}_{t\mid t-1}$ $=\boldsymbol{\theta}^{\,i}_{t-1\mid t-1},$ (90)
$\mathbf{W}^{\,i}_{t\mid t-1}$ $=\mathbf{Q}^{\,i}+\mathbf{W}^{\,i}_{t-1\mid t-1}.$ (91)
@@ 1.3 Filtering
$p(\boldsymbol{\theta}_{t}|\mathbf{x}_{0:t},\mathbf{w})$ $=\prod_{i=1}^{N}\prod_{l=1}^{L}\exp[\theta_{i,t}x^{l}_{i,t}+\sum_{j=1}^{N}\theta_{ij,t}x^{l}_{it}x^{l}_{j,t-1}-\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})]$
$\times\prod_{i=1}^{N}\exp[-\frac{1}{2}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1})^{\top}(\mathbf{W}^{i}_{t|t-1})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1})]\times const.$ (92)
This expression confirms that the filter distribution at time $t$ is a product of the individual neurons’ filter distributions, validating the assumption of independent filter distributions in constructing the one-step prediction density. The result enables independent filtering for each neuron.
We now approximate the filter density by the Gaussian distribution using Laplace’s method. Namely, we obtain the maximum a posteriori (MAP) estimate of the filter density, and use the Hessian at around the MAP estimate to obtain the approximate covariance. Using
$\boldsymbol{\theta}^{i}_{t}$ $=\left[\theta_{i,t},\theta_{i1,t},...\theta_{iN,t}\right]^{\top},$ (93)
$\mathbf{F}(x^{l}_{i,t},\mathbf{x}^{l}_{t-1})$ $=\left[x^{l}_{i,t},\,x_{i,t}^{l}x^{l}_{1,t-1},\,x^{l}_{i,t}x^{l}_{2,t-1},\,\dots,\,x^{l}_{i,t}x^{l}_{N,t-1}\right]^{\top},$ (94)
@@
$p(\boldsymbol{\theta}_{t}|\mathbf{x}_{0:t},\mathbf{w})$ $\propto\prod_{i=1}^{N}\exp\Big[\sum_{l=1}^{L}(\boldsymbol{\theta}^{i}_{t})^{\top}\mathbf{F}(x^{l}_{i,t},\mathbf{x}^{l}_{t-1})-\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})$
$\phantom{=}-\frac{1}{2}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1})^{\top}(\mathbf{W}^{i}_{t|t-1})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1})\Big],$ (95)
where $\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})$ is defined as
$\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})$ $=\log\left[1+\exp\left[(\boldsymbol{\theta}^{i}_{t})^{\top}\mathbf{F}(1,\mathbf{x}^{l}_{t-1})\right]\right].$ (96)
@@
$\boldsymbol{\theta}_{\rm MAP}=\arg\max_{\boldsymbol{\theta}_{t}}\log p(\boldsymbol{\theta}_{t}|\mathbf{x}_{0:t},\mathbf{w}).$ (97)
We obtain the MAP estimate through numerical optimization using the Newton–Raphson method. For this goal, we obtain the first and second-order derivative of the log posterior with respect to $\boldsymbol{\theta}^{i}_{t}$. The first-order derivatives with respect to $\boldsymbol{\theta}^{i}_{t}$ result in
$\frac{\partial\log p(\boldsymbol{\theta}_{t}|\mathbf{x}_{0:t},\mathbf{w})}{\partial\boldsymbol{\theta}^{i}_{t}}$ $=\sum_{l=1}^{L}\left[\mathbf{F}(x^{l}_{i,t},\mathbf{x}^{l}_{t-1})-\frac{\partial\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})}{\partial\boldsymbol{\theta}^{i}_{t}}\right]-(\mathbf{W}^{i}_{t|t-1})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1}),$ (98)
@@
$=r_{i,t}^{l}(\mathbf{x}^{l}_{t-1})\mathbf{F}(1,\mathbf{x}^{l}_{t-1}),$ (99)
where we defined the expected rate of $i$th neuron at time $t$ given the activity of the previous time step $\mathbf{x}^{l}_{t-1}$ as
$r_{i,t}^{l}(\mathbf{x}^{l}_{t-1})$ $\equiv E_{x^{l}_{i,t}|\mathbf{x}^{l}_{t-1}}[x^{l}_{i,t}]$

### v3

@@ Acknowledgments
We thank Shinji Nakaoka for his kind support of this project. This work was supported by JSPS KAKENHI Grant Number JP 20K11709, 21H05246, 24K21518, 25K03085 (H.S.).
Author Contributions
K.I. developed the algorithms, implemented the code, and performed the data analyses. H.S. conceived and supervised the project and contributed to analyses. Both authors wrote and revised the manuscript.
Competing Interests
The authors declare no competing interests.
State-space kinetic Ising model reveals task-dependent entropy flow in sparsely active nonequilibrium neuronal dynamics
Supplementary Information
@@
Graduate School of Life Science, Hokkaido University, Sapporo, Japan
@@
Supplementary Note 1 State-space kinetic Ising model
In this Supplementary Note, we provide the filtering and smoothing algorithms for the time-varying kinetic Ising model and an optimization method of its hyperparameters via the Expectation-Maximization algorithm.
@@ 1.1 Model
Let $x_{i,t}=\{0,1\}$ be an outcome of a binary random variable of neuron $i$ at time $t$ ($i=1,\ldots,N$, $t=0,\ldots,T$). In the kinetic Ising model, the activation of neuron $i$ at time $t$ independently depends on the activities of the neurons in the previous time step t1t-1. The conditional probability mass function of $x_{i,t}$ is given as
$p({x}_{i,t}|{x}_{1,t-1},\ldots,{x}_{N,t-1},\boldsymbol{\theta}^{i}_{t})=\frac{\exp\left[\theta_{i,t}x_{i,t}+\sum_{j=1}^{N}\theta_{ij,t}x_{i,t}x_{j,t-1}\right]}{1+\exp\left[\theta_{i,t}+\sum_{j=1}^{N}\theta_{ij,t}x_{j,t-1}\right]},$ (S1.1)
where $\theta_{i,t}$ is a time-dependent field parameter that determines the bias for inputs to the $i$-th neuron at time $t$, and $\theta_{ij,t}$ is a time-dependent coupling parameter from the $j$-th neuron to the $i$-th neuron at time $t$. These parameters are collectively denoted as $\boldsymbol{\theta}^{i}_{t}=(\theta_{i,t},\theta_{i1,t},\ldots\theta_{ij,t},\ldots{\theta}_{iN,t})$. Using the log normalization function,
$\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})$ $=\log\left[1+\exp\left[\theta_{i,t}+\sum_{j=1}^{N}\theta_{ij,t}x^{l}_{j,t-1}\right]\right],$ (S1.2)
@@
$p({x}_{i,t}|{x}_{1,t-1},\ldots,{x}_{N,t-1},\boldsymbol{\theta}^{i}_{t})$ $=\exp\left[\theta_{i,t}x_{i,t}+\sum_{j=1}^{N}\theta_{ij,t}x_{i,t}x_{j,t-1}-\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}_{t-1})\right].$ (S1.3)
Assuming conditional independence, the joint probability mass function that determines the probabilities of generating patterns of activity across $N$ neurons is given by
$\prod_{i=1}^{N}p({x}_{i,t}|{x}_{1,t-1},\ldots,{x}_{N,t-1},\boldsymbol{\theta}^{i}_{t}).$ (S1.4)
Typical neurophysiological experiments repeat multiple trials of measurement under the same experimental conditions. We let $x^{l}_{i,t}=\{0,1\}$ be a binary variable of the $i$-th neuron at time $t$ in the $l$-th trial ($i=1,\ldots,N$, $t=0,\ldots,T$, $l=1,\ldots,L$). We collectively denote the binary patterns of simultaneously recorded neurons at time $t$ in the $l$-th trial using a vector, $\mathbf{x}^{l}_{t}=(x^{l}_{1,t},\ldots,x^{l}_{N,t})$. Further, we denote the patterns at time $t$ from all trials by $\mathbf{x}_{t}=(\mathbf{x}^{1}_{t},\ldots,\mathbf{x}^{l}_{t},\ldots,\mathbf{x}^{L}_{t})$ and denote all the patterns up to time $t$ by $\mathbf{x}_{0:t}$. We use the same convention for the time-varying parameters, denoting them as $\boldsymbol{\theta}_{t}=(\boldsymbol{\theta}^{1}_{t},\ldots,\boldsymbol{\theta}^{i}_{t},\ldots,\boldsymbol{\theta}^{N}_{t})$ and $\boldsymbol{\theta}_{1:t}$ for their trajectories over time.
@@
$p(\mathbf{x}_{0:T}|\boldsymbol{\theta}_{1:T})$ $=\prod_{l=1}^{L}\prod_{i=1}^{N}\left[p(x^{l}_{i,0})\prod_{t=1}^{T}p({x}^{l}_{i,t}|\mathbf{x}^{l}_{t-1},\boldsymbol{\theta}^{i}_{t})\right],$ (S1.5)
where we use $p(x^{l}_{i,0})=0.5$ for data generation. We assume that the same time-dependent parameters apply across trials.
In the state-space model, the state model defines the discrete-time stochastic processes of the latent variables, which are the time-varying parameters $\boldsymbol{\theta}_{0:T}$ in our model. We use the following Gaussian model by assuming independent processes across neurons:
$p(\boldsymbol{\theta}_{0:T})$ $=\prod_{i=1}^{N}\left[p(\boldsymbol{\theta}^{i}_{0}|\boldsymbol{\mu}^{i},\boldsymbol{\Sigma}^{i})\prod_{t=1}^{T}p(\boldsymbol{\theta}^{i}_{t}|\boldsymbol{\theta}^{i}_{t-1},{Q}^{i})\right],$ (S1.6)
@@
$p(\boldsymbol{\theta}^{i}_{t}|\boldsymbol{\theta}^{i}_{t-1},\mathbf{Q}^{i})$ $=\frac{1}{\sqrt{|2\pi\mathbf{Q}^{i}|}}\exp\left[-\frac{1}{2}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})^{\top}(\mathbf{Q}^{i})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t-1})\right]$ (S1.7)
with $\mathbf{Q}^{\,i}$ being the noise covariance for the transition of the $i$-th neuron. The initial density of the $i$-th neuron $p(\boldsymbol{\theta}_{0}^{i}|\boldsymbol{\mu}^{i},\boldsymbol{\Sigma}^{i})$ is given as a Gaussian distribution with mean $\boldsymbol{\mu}^{i}$ and covariance $\boldsymbol{\Sigma}^{i}$. In practice, we used a zero vector and a unit matrix before optimization, respectively. In the followings, we denote a set of hyperparameters $\boldsymbol{\mu}^{i},\boldsymbol{\Sigma}^{i},\mathbf{Q}^{\,i}$ for $i=1,\ldots,N$ collectively by $\mathbf{w}$.
@@ 1.2 One-step prediction density
In this section, we derive the one-step prediction density $p(\boldsymbol{\theta}_{t}{|}\mathbf{x}_{0:t-1},\mathbf{w})$, using Chapman–Kolmogorov’s equation.
For $t=1$, we note that the one-step prediction is specified as a prior distribution: $p(\boldsymbol{\theta}_{1}{|}\mathbf{x}_{0},\mathbf{w})=p(\boldsymbol{\theta}_{1}{|}\mathbf{w})=\prod_{i=1}^{N}\mathcal{N}\bigl(\boldsymbol{\theta}_{1}^{i};\boldsymbol{\mu}^{i},\boldsymbol{\Sigma}^{i}\bigr)$. For $t=2,\ldots,T$, the one-step prediction density is computed via the Chapman–Kolmogorov equation:
$p(\boldsymbol{\theta}_{t}{|}\mathbf{x}_{0:t-1},\mathbf{w})$ $=\int p(\boldsymbol{\theta}_{t},\boldsymbol{\theta}_{t-1}{|}\mathbf{x}_{0:t-1},\mathbf{w})\,d\boldsymbol{\theta}_{t-1}$
$=\int p(\boldsymbol{\theta}_{t}{|}\boldsymbol{\theta}_{t-1},\mathbf{x}_{0:t-1},\mathbf{w})\;p(\boldsymbol{\theta}_{t-1}{|}\mathbf{x}_{0:t-1},\mathbf{w})\,d\boldsymbol{\theta}_{t-1},$ (S1.8)
where $p(\boldsymbol{\theta}_{t-1}{|}\mathbf{x}_{0:t-1},\mathbf{w})$ is the filter density at time t1t-1. We assume that the filter density factors into a product of individual neurons. Coupled with the factorized assumption of the state model, this leads to the factorization of the one-step prediction density:
$p(\boldsymbol{\theta}_{t}{|}\mathbf{x}_{0:t-1},\mathbf{w})$ $=\prod_{i=1}^{N}\int p(\boldsymbol{\theta}^{\,i}_{t}{|}\boldsymbol{\theta}^{\,i}_{t-1},\mathbf{w})\;p(\boldsymbol{\theta}^{\,i}_{t-1}{|}\mathbf{x}_{0:t-1},\mathbf{w})\,d\boldsymbol{\theta}^{\,i}_{t-1}.$ (S1.9)
We further assume that the filter density at time t1t-1, $p(\boldsymbol{\theta}^{\,i}_{t-1}{|}\mathbf{x}_{0:t-1},\mathbf{w})$, is approximated by a Gaussian distribution with mean $\boldsymbol{\theta}^{\,i}_{t-1{|}t-1}$ and covariance $\mathbf{W}^{\,i}_{t-1{|}t-1}$ (to be justified at the next filtering step):
$p(\boldsymbol{\theta}^{\,i}_{t-1}{|}\mathbf{x}_{0:t-1},\mathbf{w})=\mathcal{N}\bigl(\boldsymbol{\theta}^{\,i}_{t-1};\boldsymbol{\theta}^{\,i}_{t-1{|}t-1},\mathbf{W}^{\,i}_{t-1{|}t-1}\bigr).$ (S1.10)
Here the filter mean is defined as
$\boldsymbol{\theta}^{i}_{t-1|t-1}=\int p(\boldsymbol{\theta}^{i}_{t-1}|\mathbf{x}_{0:t-1})\boldsymbol{\theta}^{i}_{t-1}d\boldsymbol{\theta}^{i}_{t-1}=E_{\boldsymbol{\theta}^{i}_{t-1}|\mathbf{x}_{0:t-1}}\boldsymbol{\theta}^{i}_{t-1}.$ (S1.11)
It represents the expected value of the parameter at time t1t-1 using data up to t1t-1. The filter covariance is
$\mathbf{W}^{i}_{t-1|t-1}=E_{\boldsymbol{\theta}^{i}_{t-1}|\mathbf{x}_{0:t-1}}(\boldsymbol{\theta}^{i}_{t-1}-E_{\boldsymbol{\theta}^{i}_{t-1}|\mathbf{x}_{0:t-1}}\boldsymbol{\theta}^{i}_{t-1})(\boldsymbol{\theta}^{i}_{t-1}-E_{\boldsymbol{\theta}^{i}_{t-1}|\mathbf{x}_{0:t-1}}\boldsymbol{\theta}^{i}_{t-1})^{\top}.$ (S1.12)
Given the Gaussian transition model $p(\boldsymbol{\theta}^{\,i}_{t}{|}\boldsymbol{\theta}^{\,i}_{t-1},\mathbf{w})=\mathcal{N}(\boldsymbol{\theta}^{\,i}_{t};\boldsymbol{\theta}^{\,i}_{t-1},\mathbf{Q}^{\,i})$, the one-step prediction density $p(\boldsymbol{\theta}_{t}{|}\mathbf{x}_{0:t-1},\mathbf{w})$ becomes a Gaussian distribution. Namely, by completing the square with resect to $\boldsymbol{\theta}^{\,i}_{t}$ and calculating the integral, we obtain
$p(\boldsymbol{\theta}_{t}{|}\mathbf{x}_{0:t-1},\mathbf{w})=\prod_{i=1}^{N}\mathcal{N}\bigl(\boldsymbol{\theta}_{t};\boldsymbol{\theta}^{\,i}_{t{|}t-1},\,\mathbf{W}^{\,i}_{t{|}t-1}\bigr),$ (S1.13)
@@
$\boldsymbol{\theta}^{\,i}_{t{|}t-1}$ $=\boldsymbol{\theta}^{\,i}_{t-1{|}t-1},$ (S1.14)
$\mathbf{W}^{\,i}_{t{|}t-1}$ $=\mathbf{W}^{\,i}_{t-1{|}t-1}+\mathbf{Q}^{\,i}.$ (S1.15)
We also define $\boldsymbol{\theta}^{\,i}_{1{|}0}=\boldsymbol{\mu}^{i}$ and $\mathbf{W}^{\,i}_{t{|}0}=\boldsymbol{\Sigma}^{i}$ for the consistent notation of the one-step prediction density for $t=1,\ldots,T$ in subsequent calculations.
@@ 1.3 Filtering
$p(\boldsymbol{\theta}_{t}|\mathbf{x}_{0:t},\mathbf{w})$ $\propto\prod_{i=1}^{N}\prod_{l=1}^{L}\exp\left[\theta_{i,t}x^{l}_{i,t}+\sum_{j=1}^{N}\theta_{ij,t}x^{l}_{it}x^{l}_{j,t-1}-\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})\right]$
$\phantom{=}\cdot\prod_{i=1}^{N}\exp\left[-\frac{1}{2}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1})^{\top}(\mathbf{W}^{i}_{t|t-1})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1})\right].$ (S1.16)
This expression confirms that the filter density at time $t$ is a product of the individual neurons’ filter densities, validating the assumption of independent filter densities in constructing the one-step prediction density. The result enables independent filtering for each neuron.
We now approximate the filter density by the Gaussian distribution using Laplace’s method. Namely, we obtain the maximum a posteriori (MAP) estimate of the filter density and use the Hessian at around the MAP estimate to obtain the approximate covariance. Using
$\boldsymbol{\theta}^{i}_{t}$ $=\left[\theta_{i,t},\theta_{i1,t},...\theta_{iN,t}\right]^{\top},$ (S1.17)
$\mathbf{F}(x^{l}_{i,t},\mathbf{x}^{l}_{t-1})$ $=\left[x^{l}_{i,t},\,x_{i,t}^{l}x^{l}_{1,t-1},\,x^{l}_{i,t}x^{l}_{2,t-1},\,\dots,\,x^{l}_{i,t}x^{l}_{N,t-1}\right]^{\top},$ (S1.18)
@@
$p(\boldsymbol{\theta}_{t}|\mathbf{x}_{0:t},\mathbf{w})$ $\propto\prod_{i=1}^{N}\exp\Bigg[\sum_{l=1}^{L}(\boldsymbol{\theta}^{i}_{t})^{\top}\mathbf{F}(x^{l}_{i,t},\mathbf{x}^{l}_{t-1})-\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})$
$\phantom{=}-\frac{1}{2}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1})^{\top}(\mathbf{W}^{i}_{t|t-1})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1})\Bigg],$ (S1.19)
where $\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})$ is now given as
$\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})$ $=\log\left[1+\exp\left[(\boldsymbol{\theta}^{i}_{t})^{\top}\mathbf{F}(1,\mathbf{x}^{l}_{t-1})\right]\right].$ (S1.20)
@@
$\boldsymbol{\theta}_{\rm MAP}=\arg\max_{\boldsymbol{\theta}_{t}}\log p(\boldsymbol{\theta}_{t}|\mathbf{x}_{0:t},\mathbf{w}).$ (S1.21)
We obtain the MAP estimate through numerical optimization using the Newton-Raphson method. Notably, the MAP estimate for each neurons, $\boldsymbol{\theta}_{\rm MAP}^{i}$, can be obtained independently of the others. For this goal, we obtain the first and second-order derivatives of the log posterior with respect to $\boldsymbol{\theta}^{i}_{t}$. The first-order derivative with respect to $\boldsymbol{\theta}^{i}_{t}$ results in
$\frac{\partial\log p(\boldsymbol{\theta}_{t}|\mathbf{x}_{0:t},\mathbf{w})}{\partial\boldsymbol{\theta}^{i}_{t}}$ $=\sum_{l=1}^{L}\left[\mathbf{F}(x^{l}_{i,t},\mathbf{x}^{l}_{t-1})-\frac{\partial\psi(\boldsymbol{\theta}^{i}_{t},\mathbf{x}^{l}_{t-1})}{\partial\boldsymbol{\theta}^{i}_{t}}\right]-(\mathbf{W}^{i}_{t|t-1})^{-1}(\boldsymbol{\theta}^{i}_{t}-\boldsymbol{\theta}^{i}_{t|t-1}).$ (S1.22)
@@
$=r_{i,t}^{l}(\mathbf{x}^{l}_{t-1})\mathbf{F}(1,\mathbf{x}^{l}_{t-1}),$ (S1.23)
where we defined the expected rate of $i$-th neuron at time $t$ given the activity of the previous time step $\mathbf{x}^{l}_{t-1}$ as
$r_{i,t}^{l}(\mathbf{x}^{l}_{t-1})$ $\equiv E_{x^{l}_{i,t}|\mathbf{x}^{l}_{t-1}}x^{l}_{i,t}$